OpenStudy (anonymous):

add 8+ x-1 divided by x+4?

OpenStudy (anonymous):

\[\LARGE 8+ \frac{x-1}{x+4}\] ? or... \[\LARGE \frac{8+x-1}{x+4}\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[\LARGE 8+{x-1 \over x+4}={8(x+4)\over x+4}+{x-1\over x+4}=\] can you do it now? :)

OpenStudy (anonymous):

no it still confuses me

OpenStudy (anonymous):

ok... Don't worry, we'll do it together ;) \[\LARGE {8(x+4)\over x+4}+{x-1\over x+4}={8\cdot x+8\cdot 4 \over x+4}+{x-1\over x+4}=\] \[\LARGE {(8x+32)+(x-1)\over x+4}\] can you do it now ? :)

OpenStudy (anonymous):